OpenStudy (anonymous):

A and B are positive integers which quantity is larger? A. a/b B. a+3/b+3

myininaya (myininaya):

are a and b also integers or just A and B?

myininaya (myininaya):

and i suppose B. is suppose to be \[\frac{a+3}{b+3}\]

myininaya (myininaya):

not a+3/b+3

OpenStudy (anonymous):

I think Letter. B is a larger quantity

OpenStudy (anonymous):

a and b. A. and B are just used to seperate the answers

jimthompson5910 (jim_thompson5910):

a/b ? (a+3)/(b+3) a(b+3) ? b(a+3) ab+3a ? ab+3b 3a ? 3b a ? b So the answer depends on whether if a > b or a < b

OpenStudy (anonymous):

so then it cannot be determined?

myininaya (myininaya):

yes we need need know if a<b or if a>b so not enough info

OpenStudy (anonymous):

thanks

myininaya (myininaya):

omg i want some ice cream

OpenStudy (anonymous):

you could also subtract the fractions to see the conditions: \[\frac{a}{b} - \frac{a+3}{b+3} \iff \frac{3(a-b)}{b(b+3)}\] if a > b, then a/b is bigger, if a<b, then (a+3)\(b+3) is bigger.

OpenStudy (anonymous):

i want ice cream too =/ vanilla bean.

myininaya (myininaya):

i'm gonna get some ice cream and put some of that chocolate that hardens on it

OpenStudy (anonymous):

or you could try it with a = 1, b = 2, and then switch and make b = 1 and a = 2

OpenStudy (anonymous):

ice cream and baked bean

myininaya (myininaya):

a=icecream and b=chocalate and then switch thats what i did

jimthompson5910 (jim_thompson5910):

example, if a=1 and b=2 (so a < b), then a/b = 1/2 = 0.5 and (a+3)/(b+3) = 4/5 = 0.8 So if a=1 and b=2 (a < b), then a/b < (a+3)/(b+3) BUT.... if a=2 and b=1 (now a > b), then a/b = 2/1 = 2 and (a+3)/(b+3) = 5/4 = 1.25 So if a=2 and b=1 (a > b), then a/b > (a+3)/(b+3) See the difference?

OpenStudy (anonymous):

guess we all had the same idea

OpenStudy (anonymous):

great minds think alike :P

OpenStudy (anonymous):

no about that ice cream <.<

OpenStudy (anonymous):

ice cream for crows

OpenStudy (anonymous):

now*

myininaya (myininaya):

crows?

OpenStudy (anonymous):

i hate my color =/

myininaya (myininaya):

it looks good on you joe lol

OpenStudy (anonymous):

lolol

myininaya (myininaya):

i think the pink brings out your eyes

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

yeah you look like pepto bimsol

myininaya (myininaya):

lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

actually i do. you look more like a faded my little pony

OpenStudy (anonymous):

i didnt care for the precious color at first, but now i want it back <.<

OpenStudy (anonymous):

previous*

myininaya (myininaya):

next you will be light pink

OpenStudy (anonymous):

precious is more a propos

OpenStudy (anonymous):

lol >.<